OpenStudy (anonymous):

(1) The number of items on two grocery lists differs by 7. (2) The total number of items is 33. How many items are on each list?

OpenStudy (anonymous):

list 1 = x items list 2= (x+7) items total items on both lists = 33 can you proceed from here?

OpenStudy (anonymous):

You can set this up as a simple system of equations. If one grocery list is 'x' and one is 'y' then: x-y = 7 x+y = 33

OpenStudy (anonymous):

@qpHalcy0n you're also correct, but why use two unknowns when you can set up with one?

OpenStudy (anonymous):

i need a little more help

OpenStudy (anonymous):

You don't necessarily have to solve the system as such. Solve for either variable in one and just subtract from 33.

OpenStudy (anonymous):

@ rayleigh: list 1 +list 2 = 33 x + (x+7)=33 2x +7 =33 2x = 26 x = 13 ........list 1 x+7 = 13+7 = 20........list 2

OpenStudy (anonymous):

thank you!